# Attenuation Bias

### v2

@@ 3.2 Method of Composition (MOC): Why It Exacerbates Attenuation
Previous work in political science and elsewhere has considered how to appropriately propagate the uncertainty in estimates of latent predictors into estimated slope coefficients in regression models. Much of this work has employed an approach based on the method of composition (MOC) [45, p. 52]. In a prominent article that introduced the specific variant of MOC commonly used in political science and more broadly, [47] estimate countries’ levels of democracy, a latent trait, using Bayesian MCMC; then, for each iteration of the sampler for the latent traits, they estimate a regression predicting the risk of civil war with level of democracy and other factors, where level of democracy is set to the iteration’s sampled value of the latent traits. Many high-impact works have used MOC with the goal of appropriately accounting for the uncertainty in latent predictors [8, 13, 14, 27, 28, 30, see e.g.]. The Varieties of Democracy (V-Dem) project even provides, in addition to point estimates of their democracy measures, a series of posterior samples so that researchers can employ MOC in subsequent models [38]. [^5] Note that V-Dem’s use of MOC to combine individual measures into more general indices and account for the resulting uncertainty [35] does not necessarily suffer from the issues we identify in using MOC to fit second stage regression models with latent predictors. In a recent APSR Letter, [44] criticize previous work on the relationship between public support for democracy and regime survival, arguing that MOC should be used to incorporate the uncertainty in estimates of the latent predictors.
Although it shares a name with a general method of sampling from a distribution, MOC for latent predictors, as advocated by [47] and as commonly used in the subsequent latent traits literature, uses slightly different steps. The general MOC, used to sample from the density for some random variable $y$, first samples $x^{\prime}$ from the density for some other random variable $g(x)$, then samples $y$ from $f(y|x^{\prime})$. As applied to the latent predictor situation with a matrix of indicators $W$ from which we can estimate the latent trait $X$ according to some statistical model, MOC proceeds according to the following steps for a set of draws $t=1,\ldots,T$:
@@
MOC is typically motivated by the concern that simply using point estimates of the latent trait as a predictor may result in over-confident findings and that the uncertainty in the latent trait estimates should be taken into account in the estimation, with their uncertainty allowed to “propagate” into the regression level of the model. But as discussed above, the primary concern when estimating regression models using a predictor measured with error is attenuation bias, which results in slope coefficient estimates being pulled toward zero on average for the predictor measured with error. Furthermore, in the simple linear regression case, although standard errors can be biased either upwards or downwards, this is counteracted by the coefficient attenuation, with the overall result that $t$-statistics are biased toward zero. [^6] See [39], p. 3 for a derivation of this $t$-statistic bias. This means that overconfidence is not a problem that typically needs remediation in this situation.
@@
MOC and related approaches average over a set of estimates, each using separate draws from the posterior distribution for the latent variable given the indicator data. Each of these individual draws will typically have even more random variation than do the point estimates of the latent traits. To see this, note that the point estimates are themselves equal to the true latent traits plus some measurement error, and the individual draws from the posterior distribution for $X$ given $W$ can be thought of as being randomly distributed around the point estimates. i.e., $X^{(t)}=\hat{X}+\xi$ where $\hat{X}$ is the point estimate (e.g., the posterior mean for $X$ given $W$) and $\xi$ is an additional source of mean-zero randomness relative to just using the point estimates for $X$. MOC thus averages over a series of even further attenuated estimates as compared to using the point estimates of the latent traits themselves. The attenuation factor with MOC is now based on a measurement error variance that is the sum of the measurement error variance from the traditional case in Equation 3 and the variance of the conditional posterior for the latent traits $X$. In other words, applying MOC is analogous to increasing the measurement error from $\sigma^{2}_{u}$ to $\sigma^{2}_{u}+\sigma^{2}_{\xi}$ for each draw of the sampler, then averaging over these more strongly biased values.
Seen in this light, it is unsurprising that many applications of MOC have estimated slope coefficients on latent predictors to be even closer to zero than those from applications using ordinary regression based on the point estimates of the latent trait. In other words, MOC tends to exacerbate, rather than ameliorate, the main problem of attenuation, generally making it worse (sometimes much worse) than it would be if one had simply ignored measurement error altogether.
3.3 Full Joint Estimation: Benefits and Practical Costs
Although MOC is sometimes described as a joint modeling strategy, it is, strictly speaking, not. In the MOC implementation used with latent predictors, one first draws $X^{(t)}\sim p(X\mid W)$ and then samples $(\beta,\sigma_{\varepsilon}^{2})\sim p(\beta,\sigma_{\varepsilon}^{2}\mid Y,X^{(t)})$. Crucially, $Y$ is not conditioned on when sampling $X^{(t)}$, whereas a genuinely joint approach targets
$p(X,\beta,\theta\mid Y,W)\ \propto\ p(Y\mid X,\beta,\sigma_{\varepsilon}^{2})\,p(W\mid X,\theta)\,p(X)\,p(\beta,\sigma_{\varepsilon}^{2},\theta),$
so that the conditional for the latent trait updates with the outcome, $X\sim p(X\mid W,Y,\beta,\theta)$. Put differently, MOC integrates over $p(X\mid W)$ rather than $p(X\mid W,Y)$ and therefore ignores outcome information when updating $X$, which—given identification-induced rescaling—tends to exacerbate attenuation rather than cure it (as shown above).777[47] make this point and note that MOC effectively assumes $p(X\mid Z,Y)=p(X\mid Z)$; in our notation, their $Z$ corresponds to our indicator matrix $W$.
It is sometimes possible to estimate the measurement and outcome models jointly—for example, with Bayesian factor–outcome models or SEM/LISREL. Under correct specification, such joint estimation conditions on $Y$ when learning $X$ and avoids the extra attenuation that arises when treating a noisy, rescaled $\hat{X}$ as fixed. The costs are practical and conceptual: one must posit a full statistical model including $p(W\mid X)$ and $p(Y\mid X)$, handle scale/sign identification, and fit the joint posterior or likelihood (typically via specialized MCMC or SEM software). This can be demanding, sensitive to misspecification, and often infeasible for “black‑box” scores (e.g., PCA, additive indices, or ML/text embeddings) that lack an explicit likelihood (or one that can be tractably deployed via MCMC).
4 Corrected Estimation Using Split Indicators
@@
One approach to dealing with the specific type of attenuation that arises in the latent predictor case is to fit a standard linear regression using the latent trait estimates ($\hat{X}$) as the predictor and then correct the slope estimate based on the attenuation factor from Equation 3 above. To do this, we need a way to estimate $\sigma^{2}_{U}$ since the attenuation factor is determined by this measurement error variance. Importantly, this is measurement error variance on the scale of the true latent trait $X$, not the transformed (standardized) scale for the estimates of the latent trait that we can observe.
One way to obtain information about $\sigma^{2}_{U}$ is by having multiple measures of the latent trait $X$ for one or more observations. If these measures tend to give very similar values for each observation, then $\sigma^{2}_{U}$ is likely to be small, whereas if there are large differences between the measures, $\sigma^{2}_{U}$ is likely to be larger. In the case of estimating latent traits based on multiple indicators, we can typically obtain multiple measures by partitioning the indicators into multiple sets. For example, imagine we had 10 indicators for a latent trait $X$. We could use 5 of the indicators to obtain estimates $\hat{X}_{1}$ and use the other 5 to obtain estimates $\hat{X}_{2}$. If we split the indicators evenly, it may be reasonable to assume that the variance of the measurement errors for each of these two sets of estimates is the same. [^8] For an odd number of indicators, consider all partitions where the two sets differ by one item.
Following the notation above, we will denote two (unobserved) estimates of the latent trait on the same scale as the true latent trait itself as $\tilde{X}_{1}=X+U_{1}$ and $\tilde{X}_{2}=X+U_{2}$ where we assume that measurement errors $U_{1}$ and $U_{2}$ each have mean zero and the same variance $\sigma^{2}_{U}$ and are independent of each other and of $X$. We also define $\hat{X}_{1}$ and $\hat{X}_{2}$ as the corresponding identified estimates, which we can observe if we assume they have a mean of 0 and a variance of 1. To estimate the attenuation factor, we can first recognize that even without knowing $\tilde{X}_{1}$ or $\tilde{X}_{2}$, the correlation between $\tilde{X}_{1}$ and $\tilde{X}_{1}$ must be the same as the correlation between the observable estimates $\hat{X}_{1}$ and $\hat{X}_{2}$ since these are simply affine transformations of each other. We can then show that
@@
$=\frac{\sigma^{2}_{X}}{\sigma^{2}_{X}+\sigma^{2}_{U}}=\lambda$
@@
where, as in Equation 2 above, $\lambda$ is the attenuation factor in the standard (non-latent) bivariate linear regression case where the predictor is measured with classical error and, with a slight abuse of notation, $\sigma^{2}_{U}\equiv\textrm{Var}{(U_{1})}=\textrm{Var}{(U_{2})}$. This follows from the assumption that $U_{1}$ and $U_{2}$ are independent of each other and of $X$. In our setup, $\sigma^{2}_{X}=1$ by identification so that $\lambda=\nicefrac{{1}}{{(1+\sigma^{2}_{U})}}$.
@@
This, along with Equation 3 above, suggests using the square root of the sample correlation between the two replicate measures of $X$ in order to correct for attenuation in the estimated slope coefficient, $\beta_{X}$. [^9] In some situations, this split-half may not be appropriate. For example, if indicators are designed to tap different aspects of a latent variable rather than with internal consistency in mind (e.g., [20]), it may be preferable to consider test-retest reliability. Noting that in our latent variable predictor setup the attenuation factor for the slope coefficient is $\frac{1}{\sqrt{1+\sigma^{2}_{U}}}$ rather than the $\frac{1}{1+\sigma^{2}_{U}}$ from the standard observable variable case, we could estimate the standard linear regression coefficient using $\hat{X}_{1}$, as our predictor to obtain $\hat{\beta}^{*}_{\text{OLS},\hat{X}_{1}}$ and then correct this estimate to obtain
$\hat{\beta}^{*}_{\text{OLS},\hat{X}_{1}}=\frac{\hat{\beta}_{\text{OLS},\hat{X}_{1}}}{\sqrt{\text{Cor}(\hat{X}_{1},\hat{X}_{2})}}.$ (5)
@@
Here and below, we use the asterisk superscript to denote our corrected estimators. Note that we could have also used $\hat{X}_{2}$ as our predictor instead of $\hat{X}_{1}$ since we have no reason to privilege either of the two sets of estimates. [^10] Note that we cannot use $\hat{X}$—the point estimates of the latent trait using all of the indicators—as our predictor here since we have an estimate of the correction factor based on the amount of measurement error in $\hat{X}_{1}$ and $\hat{X}_{2}$, not in $\hat{X}$ which is estimated based on all of the available indicators rather than one partition of them. We can also average over these two estimators using the two sets of latent traits estimates, which has the benefit of reducing variability from the choice of which partition. [^11] Our averaging approach, and in particular the equivalent averaged instrumental variables (IV) approach we describe in the next section, shares commonalities with [19]. This averaging strategy also has the benefit of mitigating concerns about violations of the assumption that $\sigma^{2}_{U_{1}}=\sigma^{2}_{U_{2}}$, since if such a violation results in our correction factor being too small, then using the other split-half set of indicators would produce a correction factor that is too large and vice-versa. As we show in Appendix I, the bias induced when $\sigma^{2}_{U_{1}}\neq\sigma^{2}_{U_{2}}$ seems likely to be quite small, if not negligible, over plausible ranges of $\sigma^{2}_{U_{1}}$ and $\sigma^{2}_{U_{2}}$ when the estimator is averaged across these two values.
@@
Finally, because there are typically multiple ways to partition the indicators into two equally sized sets, we can define our corrected estimator $\hat{\beta}^{*}_{\text{OLS}}$ as the median over all such partitions, where we first average the estimators within each partition between $\hat{\beta}^{*}_{\text{OLS},\hat{X}_{1}}$ and $\hat{\beta}^{*}_{\text{OLS},\hat{X}_{2}}$. This helps guard against the influence of one or more outlying estimates for certain partitions of the indicators, producing more stable estimates.
@@
An advantage of this corrected OLS approach with a correlation-based adjustment is its relative simplicity and ease of implementation, requiring only the ability to split the indicators used to estimate the latent variable into two sets, estimate the latent trait separately using each of the sets of indicators, and calculate the correlation between the two estimates. This means that our approach can be used with virtually any method of estimating a latent trait, from simple additive or averaging approaches that do not specify a formal model to parametric models or advanced ML or AI approaches.
@@
As mentioned above, IV can also, under certain assumptions, produce consistent estimates of slope coefficients when a predictor is measured with error. In our case, we might use our two sets of estimates for the latent trait, using one to instrument for the other. But as shown in Appendix II, standard IV adjustments for measurement error will over-correct for attenuation, resulting in estimates that are inflated, i.e. biased away from zero by a factor of $\sqrt{1+\sigma^{2}_{U}}$. This suggests using the correlation between the two sets of latent traits estimates to adjust the standard IV estimate. In contrast to the bivariate OLS case considered above, in the IV case we multiply, rather than divide, our slope estimate by $\sqrt{\text{Cor}(\hat{X}_{1},\hat{X}_{2})}$. We then define our corrected IV estimator $\hat{\beta}^{*}_{\text{IV}}$ by averaging over the IV indicators using each set of estimates in a given partition as the predictor, then taking the median of those averages over all partitions, as we did above for the corrected OLS estimator.
In fact, as shown in Appendix II, the corrected IV and corrected OLS estimators are actually equivalent in the bivariate case. Despite this equivalence, one advantage of the IV approach is the ease with which it can be extended to multivariate regression. If one predictor is measured with error, we may either have other “control” variables or the predictor measured with error may itself be a control variable in which case adjusting for this may be important to prevent bias in the other coefficient(s) of interest. [^12] Note in the latter case we may not care about appropriately adjusting the IV coefficient on the predictor that is measured with error. We may even have multiple predictors measured with error, and believe that we have instruments that are suitable to address this. In both the OLS and IV cases, we employ the bootstrap to assess uncertainty, as this row-level resampling allows us to incorporate uncertainty not only in the coefficient estimate but also in the correction factor itself.

### v3

@@ 2.4 Method of Composition (MOC): Why It Exacerbates Attenuation
A different—and in applied latent-trait work, common—approach is the method of composition (MOC), which aims to propagate uncertainty from the measurement stage, where $X$ is estimated based on a set of observed indicators $W$, into the second-stage regression. In the MOC implementation used with latent predictors, one first draws $X^{(t)}\sim p(X\mid W)$ and then samples $(\beta,\sigma_{\varepsilon}^{2})\sim p(\beta,\sigma_{\varepsilon}^{2}\mid Y,X^{(t)})$. Crucially, $Y$ is not conditioned on when sampling $X^{(t)}$ despite it being informative about $X$ whenever $\beta\neq 0$.444[undefao] note that MOC effectively assumes $p(X\mid W,Y)=p(X\mid W)$ in our notation.
Many high-impact works have used MOC with the goal of appropriately accounting for the uncertainty in latent predictors [undefg, undefm, undefn, undefw, undefx, undefz, undefal, e.g.]. The Varieties of Democracy (V-Dem) project even provides, in addition to point estimates of their democracy measures, a series of posterior samples so that researchers can employ MOC in subsequent analyses [undefaf]. [^5] Our critique concerns using MOC to fit second-stage regression models with latent predictors. V-Dem’s use of MOC to combine individual measures into more general indices and account for the resulting uncertainty [undefac] does not necessarily suffer from the same issues.
As applied to the latent predictor situation with a matrix of indicators $W$ from which we can estimate the latent trait $X$ according to some statistical model, MOC proceeds according to the following steps for a set of draws $t=1,\ldots,T$:
@@
This two-stage sampling-and-refitting procedure is typically justified as a way to “propagate” measurement-stage uncertainty forward into the regression and avoid overstating precision. However, when the regressor is measured with error, the dominant first‑order problem is attenuation bias: measurement error pulls slope estimates toward zero.
@@
Indeed, MOC averages over a set of estimates, each using separate draws from the posterior distribution for the latent variable given the indicator data. Each of these individual draws will typically have even more random variation than do the point estimates of the latent traits. To see this, note that the point estimates are themselves equal to the true latent traits plus some measurement error, and the individual draws from the posterior distribution for $X$ given $W$ can be thought of as being randomly distributed around the point estimates. i.e., $X^{(t)}=\hat{X}+\xi$ where $\hat{X}$ is the point estimate (e.g., the posterior mean for $X$ given $W$) and $\xi$ is an additional source of mean-zero randomness relative to just using the point estimates for $X$.
3 A Correlation-Corrected Estimator Using Split Indicators
@@
Taken together—naïve IV’s tendency to over-correct once identification rescaling is imposed, and MOC’s tendency to add noise and worsen attenuation—these limitations motivate a different tack: estimate the latent-scale measurement error and then undo the identification-aware attenuation implied by Equation 5. To do this, we need a way to estimate $\sigma^{2}_{U}$, since the attenuation factor is determined by this measurement-error variance. Importantly, this is measurement error variance on the scale of the true latent trait $X$, not the transformed (standardized) scale for the estimates of the latent trait that we can observe.
One way to learn about $\sigma_{U}^{2}$—and therefore about the latent-predictor attenuation factor in Equation 5—is to obtain (at least) two independent measurements of the same latent trait $X$ on the latent scale (i.e., before the identification/standardization step that produces $\hat{X}$). This “two-measures” idea is the same basic ingredient that motivated the IV discussion above. Intuitively, if two measurements of $X$ give very similar values for each observation, then the measurement error variance $\sigma_{U}^{2}$ must be small; if the two measurements often disagree substantially, then $\sigma_{U}^{2}$ must be larger.
In indicator-based latent-trait settings, we can usually construct such repeated measures by partitioning the indicators into two disjoint sets and estimating the trait separately within each set. For example, with 10 indicators for a latent trait $X$, we could use 5 indicators to obtain estimates $\hat{X}_{1}$ and the other 5 to obtain estimates $\hat{X}_{2}$. If we split the indicators evenly, it may be reasonable to assume that the two resulting estimates have the same measurement-error variance. [^6] For an odd number of indicators, consider all partitions where the two sets differ by one item. Following the notation above, let $\tilde{X}_{1}$ and $\tilde{X}_{2}$ denote the two unobserved (unidentified) measurements on the same scale as the true latent trait:
$\tilde{X}_{1}=X+U_{1}\qquad\text{and}\qquad\tilde{X}_{2}=X+U_{2},$
where we assume $U_{1}$ and $U_{2}$ each have mean zero and the same variance $\sigma_{U}^{2}$, and are independent of each other and of XX. [^7] This independence assumption is similar to the conditional independence assumption commonly used in latent traits measurement strategies including many item response models. Finally, let $\hat{X}_{1}$ and $\hat{X}_{2}$ denote the corresponding identified versions of these split-indicator estimates—i.e., the quantities we actually observe after imposing the usual mean-zero, unit-variance normalization (consistent with Equation 4).
Although we do not observe $\tilde{X}_{1}$ or $\tilde{X}_{2}$ directly, their correlation is identical to $\operatorname{Cor}(\hat{X}_{1},\hat{X}_{2})$ because $\hat{X}_{1}$ and $\hat{X}_{2}$ are affine (mean/variance) transformations of $\tilde{X}_{1}$ and $\tilde{X}_{2}$. The idea of then using the correlation between multiple sets of estimates to correct for measurement error has been around since at least [undefaj]. Here we can then show that
@@
$=\frac{\sigma^{2}_{X}}{\sigma^{2}_{X}+\sigma^{2}_{U}}=\lambda,$
@@
where, as in Equation 3 above, $\lambda$ is the attenuation factor in the standard (non-latent) bivariate linear regression case where the predictor is measured with classical error and, with a slight abuse of notation, $\sigma^{2}_{U}\equiv\textrm{V}{(U_{1})}=\textrm{V}{(U_{2})}$. This follows from the assumption that $U_{1}$ and $U_{2}$ are independent of each other and of $X$. In our setup, $\sigma^{2}_{X}=1$ by identification so that $\lambda=\nicefrac{{1}}{{(1+\sigma^{2}_{U})}}$.
@@
In our latent variable predictor setup, the attenuation factor for the slope coefficient is $\frac{1}{\sqrt{1+\sigma^{2}_{U}}}$ (see Equation 5). In the standard observable variable case, it is $\frac{1}{1+\sigma^{2}_{U}}$ (see Equation 3). This suggests using the square root of the sample correlation between the two separate estimates of $X$ to correct attenuation in the estimated slope coefficient, $\beta_{X}$. [^8] In some situations, this split-half may not be appropriate. For example, if indicators are designed to tap different aspects of a latent variable rather than with internal consistency in mind (e.g., [undefs]), it may be preferable to consider test-retest reliability. To do this, we first estimate the standard linear regression coefficient using $\hat{X}_{1}$, as our predictor to obtain $\hat{\beta}_{\hat{X}_{1}}$ and then correct this estimate to obtain
$\hat{\beta}^{*}_{\hat{X}_{1}}=\frac{\hat{\beta}_{\hat{X}_{1}}}{\sqrt{\text{Cor}(\hat{X}_{1},\hat{X}_{2})}}.$ (7)
@@
Here and below, we use the asterisk superscript to denote our correlation-corrected estimators. Because we have no reason to privilege either split-half estimate, we compute both $\hat{\beta}^{*}_{\hat{X}_{1}}$ and $\hat{\beta}^{*}_{\hat{X}_{2}}$ and average them within a partition. [^9] We do not use $\hat{X}$—the point estimates of the latent trait using all of the indicators—because our correction factor is estimated from the split-half reliability of $\hat{X}_{1}$ and $\hat{X}_{2}$, not the full-indicator score; applying the same correction to $\hat{X}$ would require additional assumptions or estimation. This averaging reduces variability from the choice of partition. [^10] Our averaging approach, and in particular the equivalent averaged instrumental variables (IV) approach we describe in the next section, shares commonalities with [undefr]. Averaging across the two split-half estimates also makes the method less sensitive to unequal error variances. If one split yields a correction that is too small, the other tends to yield a correction that is too large. As we show in Appendix I, the bias induced when $\sigma^{2}_{U_{1}}\neq\sigma^{2}_{U_{2}}$ seems likely to be quite small, if not negligible, over plausible ranges of $\sigma^{2}_{U_{1}}$ and $\sigma^{2}_{U_{2}}$ when the estimator is averaged across these two values.
@@
Finally, because there are typically multiple ways to partition the indicators into two equally sized sets, we can define our corrected estimator $\hat{\beta}^{*}$ as the median over all such partitions, where we first average the estimators within each partition between $\hat{\beta}^{*}_{\hat{X}_{1}}$ and $\hat{\beta}^{*}_{\hat{X}_{2}}$. This helps guard against the influence of one or more outlying estimates for certain partitions of the indicators, producing more stable estimates. We employ the bootstrap to assess uncertainty, as this row-level resampling allows us to incorporate uncertainty not only in the coefficient estimate but also in the correction factor itself.
@@
An advantage of this corrected OLS approach with a correlation-based adjustment is its relative simplicity and ease of implementation. It requires only splitting the indicators used to estimate the latent variable into two sets, estimating the latent trait separately using each set of indicators, and calculating the correlation between the two estimates. This means that our approach can be used with virtually any method of estimating a latent trait, from simple additive or averaging approaches that do not specify a formal model to parametric models or advanced ML or AI approaches. In contrast, a full joint estimation of the latent traits and the regression model will provide the most principled results, but is often quite cumbersome and in cases where the latent trait estimation does not specify a full statistical model this joint approach may not be possible. As we illustrate below, our approach can produce results very similar to full joint estimation.
@@
The same split-indicator idea can also be cast in an IV framework, which is especially useful once we move beyond the bivariate case to multivariate regressions where one or more covariates are measured with error. In our case, we might use our two sets of estimates for the latent trait, using one to instrument for the other. But as shown above standard IV adjustments for measurement error will over-correct for attenuation, resulting in estimates that are inflated, i.e., biased away from zero by a factor of $\sqrt{1+\sigma^{2}_{U}}$. This suggests using the correlation between the two sets of latent trait estimates to adjust the standard IV estimate. In contrast to the bivariate OLS case considered above, in the IV case we multiply, rather than divide, our slope estimate by $\sqrt{\text{Cor}(\hat{X}_{1},\hat{X}_{2})}$. In fact, as shown in Appendix II, the correlation-corrected IV and correlation-corrected OLS estimators are equivalent in the bivariate case. Despite this equivalence, one advantage of the IV approach is the ease with which it can be extended to multivariate regression. In multivariate regressions, the error-prone variable may be the main predictor or merely a “control”; in either case, failing to account for its measurement error can bias the estimated coefficients on the other covariates of interest. We may even have multiple predictors measured with error, and believe that we have instruments that are suitable to address this.